# GVE-Louvain: Fast Louvain Algorithm for Community Detection in Shared Memory Setting

### 08 Dec 2023 v1

@@ 4.2.1. Main step of GVE-Louvain
The main step of GVE-Louvain (louvian() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, and the aggregation phase. Here, the louvain() function takes the input graph $G$, and returns the community membership $C$ for each vertex. In line 15, we first initializes the community membership $C$ of each vertex in $G$, and perform passes of the Louvain algorithm, limited to $MAX\_PASSES$ (lines 16-26). In each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 17); and mark all vertices as unprocessed (line 18).
@@ 5.2. Comparing Performance of GVE-Louvain
Figure 8. Speedup of GVE-Louvain with increasing number of threads (in multiples of 2).
@@ 5.4. Strong Scaling of GVE-Louvain
Finally, we measure the strong scaling performance of GVE-Louvain. To this end, we adjust the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the time taken for finding communities with GVE-Louvain (five times for averaging). The results are shown in Figure 8. With 32 threads, GVE-Louvain obtains a 10.4×10.4$\times$10.4 × speedup compared to running with a single thread, i.e., its performance increases by 1.6×1.6$\times$1.6 × for every doubling of threads. Scaling is limited due to the various sequential steps/phases in the algorithm. At 64 threads, GVE-Louvain is impacted by NUMA effects, and offer speedups of only 11.4×11.4$\times$11.4 ×.

### 11 Dec 2023 v2

@@ 4.2.1. Main step of GVE-Louvain
The main step of GVE-Louvain (louvian() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, and the aggregation phase. Here, the louvain() function takes the input graph $G$, and returns the community membership $C$ for each vertex. In line 15, we first initialize the community membership $C$ of each vertex in $G$, and perform passes of the Louvain algorithm, limited to $MAX\_PASSES$ (lines 16-26). In each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 17); and mark all vertices as unprocessed (line 18).
@@ 5.2. Comparing Performance of GVE-Louvain
Figure 8. Overall speedup of GVE-Louvain, and its various phases (local-moving, aggregation, others), with increasing number of threads (in multiples of 2).
@@ 5.4. Strong Scaling of GVE-Louvain
Finally, we measure the strong scaling performance of GVE-Louvain. To this end, we adjust the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the overall time taken for finding communities with GVE-Louvain, as well as its phase splits (local-moving, aggregation, others), five times for averaging. The results are shown in Figure 8. With 32 threads, GVE-Louvain obtains an average speedup of 10.4×10.4$\times$10.4 × compared to running with a single thread, i.e., its performance increases by 1.6×1.6$\times$1.6 × for every doubling of threads. Scaling is limited due to the various sequential steps/phases in the algorithm. At 64 threads, GVE-Louvain is impacted by NUMA effects, and offers speedup of only 11.4×11.4$\times$11.4 ×.